# Conjugate generation of sporadic almost simple groups

### v1

@@ 1. Introduction
Let $S$ be a nonabelian simple group (which we identify with $\operatorname{Inn}S$) and let $x\in\operatorname{Aut}S$ be a nonidentity automorphism (possibly, inner). R. Guralnick and J. Saxl [26] introduced the notation $\alpha(x)=\alpha^{\phantom{S}}_{S}(x)$ for the minimum number of elements conjugate to $x$ in $G=\langle x,S\rangle$ that generate $G$. Thus, $\alpha^{\phantom{S}}_{S}(x)$ has a constant value on every conjugacy class $x^{S}=x^{G}$ and thus may be viewed as a numeric invariant of the class. Even before [26] appeared, J. Moori [34, 35, 36] called $\alpha^{\phantom{S}}_{S}(x)$ the rank [^1] We note that Moori and some other authors, e. g. in [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 28, 34, 35, 36], use the term “rank” in the case when $x$ induces on $S$ an inner automorphism. of class $x^{S}$.
@@
A significant number of these papers go back to the pioneering work by Moori [34, 35, 36] who proved as early as 1993–1994 [34, 35] that if $S=F_{22}$ and ${x\in S}$ is a $3$-transposition then $\alpha^{\phantom{S}}_{S}(x)\in\{5,6\}$. Subsequently, J. Hall and L. Soicher [27, Theorems (1.1)–(1.3)] classified the groups of $3$-transpositions that may be generated by at most five $3$-transpositions. In particular, it was shown that the Fischer groups $Fi_{22}$ and $Fi_{23}$ cannot be generated by five $3$-transpositions, which confirmed Moori’s conjecture [35] that $\alpha^{\phantom{S}}_{S}(x)=6$ when $S=Fi_{22}$ and $x$ is a $3$-transposition. With his paper [36], Moori initiated a systematic calculation of the ranks of inner conjugacy classes of sporadic simple groups. In a series of subsequent articles by J. Moori, F. Ali, A. B. M. Basheer, M. A. F. Ibrahim, M. A. Al-Kadhi, et al. [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13], the ranks of inner conjugacy classes were found for many sporadic groups and the Tits group.
@@
      $(S,x)\in\{(J_{2},3A),(HS,4A),(McL,3A),(Ly,3A),(Co_{1},3A),(F_{22},3A),\allowbreak(Fi_{22},3B)\}$ and $\alpha^{\phantom{S}}_{S}(x)=3$,
@@
In our recent paper [41], it was shown that, for $x\in\operatorname{Aut}S\setminus S$, we either have $\alpha^{\phantom{S}}_{S}(x)\leqslant 4$, or $(S,x)=({{Fi_{24}}^{\prime}},2C)$ and $\alpha^{\phantom{S}}_{S}(x)=5$. The peculiarity of class $2C$ in $\operatorname{Aut}({Fi_{24}}^{\prime})=Fi_{24}$ is that it is a class of $3$-transpositions. The fact that $Fi_{24}$ is generated by five $3$-transpositions was established by S. Norton [37], and that $Fi_{24}$ cannot be generated by four $3$-transpositions was proved in [27, Theorem (1.1)].
@@ 2. Preliminaries
see [29, Exercise (3.9), p. 45]. For our purposes, the value of $\mathrm{m}(a,b,c)$ can thus be determined either from the Atlas [16] or using the computer algebra system GAP [19] which has the relevant built-in function ClassMultiplicationCoefficient().
@@
$(i)$ The existence of $s$ as stated is proven in [25, Proposition 6.2]. This implies that if $x\in nX$ then $S\leqslant\langle x,s\rangle$ for some $s\in S$ of order $p$. Since $\langle s\rangle$ is a Sylow $p$-subgroup of $S$, we may assume that $s\in pA$. Using [19] we check that $\operatorname{m}(nX,nX,pA)>0$ in all cases under consideration, see [42] for a detailed GAP code. This implies that $s=x_{1}x_{2}$ for some $x_{1},x_{2}\in nX$. Therefore, $S\leqslant\langle x,x_{1},x_{2}\rangle$ and so $\alpha^{\phantom{S}}_{S}(x)\leqslant 3$.
@@ 3. Proof of main results
The finite simple groups generated by a conjugacy class $D$ of elements of order $3$ such that every two members of $D$ generates one of the groups (2) were classified in  [43]. The structure of subgroups generated by three elements in $D$ was clarified in [43, Section 1], whence it follows that $Suz$ cannot be such a subgroup. In particular, $\alpha^{\phantom{S}}_{S}(x)>3$ and so $\alpha^{\phantom{S}}_{S}(x)=4$ by [17, Theorem 3.1].
@@
$\bullet$ Let $(S,nX)=({Fi_{24}}^{\prime},2D)$. In this case, we have $\operatorname{m}(2D,2D,33A)=33$ and $\operatorname{m}(2D,33A,46A)=172322171820$. Thus, there are three elements in $2D$ that generate a subgroup $H$ not being included in $S$, having order divisible by $11$ and $23$, and containing an element of order $33$. Every maximal subgroup of $G$ of order divisible by both $11$ and $23$ that is not included in $S$ is isomorphic to either $Fi_{23}\times 2$ or ${2^{12}.M_{24}}$, see [32, Theorem 1.1 and Table 1.1]. But these groups contain no elements of order $33$. Again, $H=G$ and $\alpha^{\phantom{S}}_{S}(x)=3$.
@@
$\bullet$ Let $(S,nX)$ be one of $(M_{22},2B)$, $(HS,2C)$, or $(Fi_{22},2D)$. In this case, as we mentioned in the introduction, $G=\operatorname{Aut}S$ is a $4$-transposition group with respect to the class $D=nX$. This can be verified by checking that $\operatorname{m}(x,x,y)\neq 0$ only when $y$ has order at most $4$. By [41, Theorem 2], we know that $3\leqslant\alpha^{\phantom{S}}_{S}(x)\leqslant 4$. However, $3$-generated $4$-transposition groups have been classified, see [30, Sections 3.1, 3.2] and [31, Theorem 3.4.2]. This classification implies, in particular, that $S$ cannot be a composition factor of a $3$-generated $4$-transposition group as the only possible such factor is $L_{2}(7)$. This shows that $\alpha^{\phantom{S}}_{S}(x)=4$.
@@ 4. Final remarks
We also mention some problems that remain open for sporadic groups. Theorem 1 gives no information on the value of $\alpha^{\phantom{S}}_{S}(x)$ in the case where $S$ is the Monster. Recall that, in this case, [17, Theorem 3.1] implies $\alpha^{\phantom{S}}_{S}(x)\in\{2,3\}$ if $|x|>2$ and $\alpha^{\phantom{S}}_{S}(x)\in\{3,4\}$ if $|x|=2$. The precise value of $\alpha^{\phantom{S}}_{S}(x)$ is also unknown when $S$ is sporadic and $x\in\operatorname{Aut}(S)\setminus S$ has composite order. In many such cases, there is a power $y\neq 1$ of $x$ such that $y\in S$ and $\alpha^{\phantom{S}}_{S}(y)=2$. Then we clearly have $\alpha^{\phantom{S}}_{S}(x)=2$. Despite these remarks, many open cases remain whose treatment is beyond the scope of this short paper.
@@
This is a preprint of the paper of the same title accepted for publication in the journal ‘‘Siberian Advances in Mathematics’’ (© Sobolev Institute of Mathematics of SB RAS, 2025).

### v2

@@ 1. Introduction
Let $S$ be a nonabelian simple group (which we identify with $\operatorname{Inn}S$) and let $x\in\operatorname{Aut}S$ be a nonidentity automorphism (possibly, inner). R. Guralnick and J. Saxl [27] introduced the notation $\alpha(x)=\alpha^{\phantom{S}}_{S}(x)$ for the minimum number of elements conjugate to $x$ in $G=\langle x,S\rangle$ that generate $G$. Thus, $\alpha^{\phantom{S}}_{S}(x)$ has a constant value on every conjugacy class $x^{S}=x^{G}$ and so may be viewed as a numeric invariant of the class. Even before [27] appeared, J. Moori [35, 36, 37] called $\alpha^{\phantom{S}}_{S}(x)$ the rank [^1] We note that Moori and some other authors, e. g. in [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 29, 35, 36, 37], use the term “rank” in the case when $x$ induces on $S$ an inner automorphism. of class $x^{S}$.
@@
A significant number of these papers go back to the pioneering work by Moori [35, 36, 37] who proved as early as 1993–1994 [35, 36] that if $S=F_{22}$ and ${x\in S}$ is a $3$-transposition then $\alpha^{\phantom{S}}_{S}(x)\in\{5,6\}$. Subsequently, J. Hall and L. Soicher [28, Theorems (1.1)–(1.3)] classified the groups of $3$-transpositions that may be generated by at most five $3$-transpositions. In particular, it was shown that the Fischer groups $Fi_{22}$ and $Fi_{23}$ cannot be generated by five $3$-transpositions, which confirmed Moori’s conjecture [36] that $\alpha^{\phantom{S}}_{S}(x)=6$ when $S=Fi_{22}$ and $x$ is a $3$-transposition. With his paper [37], Moori initiated a systematic calculation of the ranks of inner conjugacy classes of sporadic simple groups. In a series of subsequent articles by J. Moori, F. Ali, A. B. M. Basheer, M. A. F. Ibrahim, M. A. Al-Kadhi, et al. [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13], the ranks of inner conjugacy classes were found for many sporadic groups and the Tits group.
@@
      $(S,x)\in\{(J_{2},3A)$, $(HS,4A)$, $(McL,3A)$, $(Ly,3A)$, $(Co_{1},3A)$,
      $(F_{22},3A)$, $(Fi_{22},3B)$, $(Fi_{23},3A)$, $(Fi_{23},3B)$, $({Fi_{24}}^{\prime},3A)$, $({Fi_{24}}^{\prime},3B)\}$ and $\alpha^{\phantom{S}}_{S}(x)=3$,
@@
In our recent paper [42], it was shown that, for $x\in\operatorname{Aut}S\setminus S$, we either have $\alpha^{\phantom{S}}_{S}(x)\leqslant 4$, or $(S,x)=({{Fi_{24}}^{\prime}},2C)$ and $\alpha^{\phantom{S}}_{S}(x)=5$. The peculiarity of class $2C$ in $\operatorname{Aut}({Fi_{24}}^{\prime})=Fi_{24}$ is that it is a class of $3$-transpositions. The fact that $Fi_{24}$ is generated by five $3$-transpositions was established by S. Norton [38], and that $Fi_{24}$ cannot be generated by four $3$-transpositions was proved in [28, Theorem (1.1)].
@@
Remark. After the original version of this preprint had been published, we received a letter from Prof. J. Müller dated May 15, 2025 mentioning the paper [18] of 2019, where the values of $\alpha^{\phantom{S}}_{S}(x)$ were determined independently using other methods for all sporadic groups $S$ and all nonidentity $x\in\operatorname{Aut}S$. Theorems 1 and 2 of the present paper are particular cases of the results in [18] whose existence we had had no knowledge of. We are thankful to Prof. Müller for this information and for pointing out an inaccurate citation of [17] in our original statement of Theorem 1.
@@ 2. Preliminaries
see [30, Exercise (3.9), p. 45]. For our purposes, the value of $\mathrm{m}(a,b,c)$ can thus be determined either from the Atlas [16] or using the computer algebra system GAP [20] which has the relevant built-in function ClassMultiplicationCoefficient().
@@
$(i)$ The existence of $s$ as stated is proven in [26, Proposition 6.2]. This implies that if $x\in nX$ then $S\leqslant\langle x,s\rangle$ for some $s\in S$ of order $p$. Since $\langle s\rangle$ is a Sylow $p$-subgroup of $S$, we may assume that $s\in pA$. Using [20] we check that $\operatorname{m}(nX,nX,pA)>0$ in all cases under consideration, see [43] for a detailed GAP code. This implies that $s=x_{1}x_{2}$ for some $x_{1},x_{2}\in nX$. Therefore, $S\leqslant\langle x,x_{1},x_{2}\rangle$ and so $\alpha^{\phantom{S}}_{S}(x)\leqslant 3$.
@@ 3. Proof of main results
The finite simple groups generated by a conjugacy class $D$ of elements of order $3$ such that every two members of $D$ generates one of the groups (2) were classified in  [44]. The structure of subgroups generated by three elements in $D$ was clarified in [44, Section 1], whence it follows that $Suz$ cannot be such a subgroup. In particular, $\alpha^{\phantom{S}}_{S}(x)>3$ and so $\alpha^{\phantom{S}}_{S}(x)=4$ by [17, Theorem 3.1].
@@
$\bullet$ Let $(S,nX)=({Fi_{24}}^{\prime},2D)$. In this case, we have $\operatorname{m}(2D,2D,33A)=33$ and $\operatorname{m}(2D,33A,46A)=172322171820$. Thus, there are three elements in $2D$ that generate a subgroup $H$ not being included in $S$, having order divisible by $11$ and $23$, and containing an element of order $33$. Every maximal subgroup of $G$ of order divisible by both $11$ and $23$ that is not included in $S$ is isomorphic to either $Fi_{23}\times 2$ or ${2^{12}.M_{24}}$, see [33, Theorem 1.1 and Table 1.1]. But these groups contain no elements of order $33$. Again, $H=G$ and $\alpha^{\phantom{S}}_{S}(x)=3$.
@@
$\bullet$ Let $(S,nX)$ be one of $(M_{22},2B)$, $(HS,2C)$, or $(Fi_{22},2D)$. In this case, as we mentioned in the introduction, $G=\operatorname{Aut}S$ is a $4$-transposition group with respect to the class $D=nX$. This can be verified by checking that $\operatorname{m}(x,x,y)\neq 0$ only when $y$ has order at most $4$. By [42, Theorem 2], we know that $3\leqslant\alpha^{\phantom{S}}_{S}(x)\leqslant 4$. However, $3$-generated $4$-transposition groups have been classified, see [31, Sections 3.1, 3.2] and [32, Theorem 3.4.2]. This classification implies, in particular, that $S$ cannot be a composition factor of a $3$-generated $4$-transposition group as the only possible such factor is $L_{2}(7)$. This shows that $\alpha^{\phantom{S}}_{S}(x)=4$.
@@ 4. Final remarks
This is a preprint of the paper of the same title accepted for publication in the journal ‘‘Siberian Electronic Mathematical Reports’’ (© Sobolev Institute of Mathematics of SB RAS, 2025).